OpenStudy (anonymous):

here Choose the point-slope form of the equation below that represents the line that passes through the points (−6, 4) and (2, 0). y − 4 = −one half(x + 6) y − 4 = 2(x + 6) y + 6 = −one half(x − 4) y + 6 = 2(x − 4) Question 2 (Multiple Choice Worth 2 points) Given the equation y − 3 = one half(x + 6) in point-slope form, identify the equation of the same line in standard form. x − 2y = −12 y = one halfx y = one halfx + 6 y = one halfx + 9 Question 3 (Multiple Choice Worth 2 points) Given the equation y − 4 = three fourths(x + 8) in point-slope form, identif

OpenStudy (anonymous):

the other guy left idk why

OpenStudy (calculusfunctions):

OK what do you need to do first?

OpenStudy (calculusfunctions):

First you have to find the slope. Can you do that?

OpenStudy (anonymous):

i have no clue mabe put the variables on one side

OpenStudy (calculusfunctions):

I'll be right back.

OpenStudy (anonymous):

1. Find the slope using the formula y2-y1/x2-x1 with the given values of x and y. 2. Organize the equation into form y = mx + c after you have the x and y-intercepts. 3. The line equation you get will be your answer.

OpenStudy (calculusfunctions):

slope is the rtio of the change in y to the change in x.

OpenStudy (calculusfunctions):

OK @mikala1 what are the two points you're given?

OpenStudy (anonymous):

6,4 and 2,0

OpenStudy (calculusfunctions):

(-6, 4) and (2, 0) right?

OpenStudy (anonymous):

So those are the values you need to find the slope.

OpenStudy (calculusfunctions):

Yes so then if you compare the first y to the second y, what change has occurred? Has it increased or decreased and by how much?

OpenStudy (anonymous):

it have derease

OpenStudy (calculusfunctions):

By how much?

OpenStudy (anonymous):

2

OpenStudy (calculusfunctions):

NO! Look again.

OpenStudy (anonymous):

what one are we talking about

OpenStudy (anonymous):

5

OpenStudy (calculusfunctions):

If you compare the first y to the second y, what change has occurred? Has it increased or decreased and by how much?

OpenStudy (anonymous):

what number is the second y

OpenStudy (calculusfunctions):

Every point is (x, y).

OpenStudy (hba):

\[m=\frac{ y2-y1 }{ x2-x1}\]

OpenStudy (anonymous):

i do not know because im not good with negitive numbers

OpenStudy (calculusfunctions):

The first point is (6, 4) and the second point is (2, 0). So what are the two y values?

OpenStudy (anonymous):

4 and 0 right

OpenStudy (hba):

First Of All You Need To Find The Slope (x1,y1)=(-6, 4) And (x2,y2)= (2, 0).

OpenStudy (calculusfunctions):

@hba I know the formula. Thank you very much but I am trying to teach her so would you kindly allow me to continue? Thank you!!!

OpenStudy (hba):

Oh I See,No Problem @calculusfunctions

OpenStudy (calculusfunctions):

Yes @mikala1 the two y's are 4 and 0 so what change has occurred from the first y to the second y?

OpenStudy (anonymous):

4

OpenStudy (calculusfunctions):

Is that an increase or a decrease of 4?

OpenStudy (anonymous):

decr4ease

OpenStudy (calculusfunctions):

Right so then we say that the change in y is -4. Do you understand so far?

OpenStudy (anonymous):

some what but keep going

OpenStudy (calculusfunctions):

Now what are the two x coordinates?

OpenStudy (anonymous):

is d the ansure just asking

OpenStudy (anonymous):

-6 and 2

OpenStudy (calculusfunctions):

So then what change has occurred from the first x to the second x?

OpenStudy (anonymous):

5

OpenStudy (calculusfunctions):

NO! try again.

OpenStudy (anonymous):

9

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

any one homee

OpenStudy (calculusfunctions):

NO! it's 8. Check again.

OpenStudy (anonymous):

im sorry i was one off i am not good with - numbers

OpenStudy (calculusfunctions):

So then the change in y is -4 and the change in x is 8, correct?

OpenStudy (anonymous):

i think so sure if you say yes

OpenStudy (calculusfunctions):

The slope of a line m = (the change in y)/(the change in x) which is what @hba was telling you by rudely interrupting.

OpenStudy (calculusfunctions):

So now @mikala1 can you tell me what the slope would be?

OpenStudy (anonymous):

umm d

OpenStudy (anonymous):

hello any one there i only have 10 min left plz

OpenStudy (calculusfunctions):

Yes @mikala1 so the slope is m = -4/8 = -1/2. Do you understand?

OpenStudy (anonymous):

yes ssssure can we move on now plz im running out of ttime

OpenStudy (anonymous):

i only have 5 min left sorry to rush u

OpenStudy (anonymous):

the seond on is b right

OpenStudy (anonymous):

is any one here still

OpenStudy (calculusfunctions):

OK now the equation of a line is y = mx + b so now we know that m = -1/2. So now we use one of the points, let's say (-6, 4) and solve for b. I'll just show you since you're running out of time. x = -6, y = 4, m = -1/2 thus 4 = (-1/2)(-6) + b 4 = 3 + b b = 1 ∴ y = (-1/2)x + 1 Which is the same as answer a). Understand? Sorry I didn't want to just end up giving you the answer but you're saying you're in a rush plus this @hba guy shows no respect for anyone so it is not making my job easier.

OpenStudy (anonymous):

im sorry its just my math class is over in 2 min and i neede help thank you so much for your time

OpenStudy (anonymous):

that is all for now thank you

OpenStudy (anonymous):

@starlord6200

OpenStudy (anonymous):

why you tag me

OpenStudy (anonymous):

sorry just tried to view ur profile but it wouldnt let me unless I posted it